at work one day erica received 36 packages. Speedy delivery delivered three times as many as ralphs express, While ralphs express delivered four more packages the sendquick package service. how many packages did each service deliver to erica?
Answer by richwmiller(17219)

s=3r
r=q+4
q=r-4
r+s+q=36
r+3r+r-4=36
5r-4=36
5r=40
r=8
q = 4, r = 8, s = 24